Given a right triangle, use the Pythagorean Theorem to find "a" (one of the legs) when c = 85 and b = 53. (Round your answer to the nearest tenth.)
Answer by rfer(16322)

85^2-53^2=
7225-2809=4416
sqrt 4416=66.5